OpenStudy (anonymous):

x-3/2=-5 A.-16 B.-4 C.-7 D.-13

OpenStudy (anonymous):

- - Solve: Simplify both sides of the equation. \[x+\frac{ -3 }{ 2 }=-5\] Add 3/2 to both sides. \[x+\frac{ -3 }{ 2 }+\frac{ 3 }{ 2 }=-5+\frac{ 3 }{ 2 }\] Answer: \[x=\frac{ -7 }{ 2}\]

OpenStudy (anonymous):

Medal if I helped?

OpenStudy (anonymous):

Thank you